# H-FA: A Hybrid Floating-Point and Logarithmic Approach to Hardware Accelerated FlashAttention

### v1

@@ I Introduction
Modern ML and AI systems rely on deep learning models with billions of parameters, reaching human-level performance in tasks like image recognition [1] and natural language processing [2]. Central to this progress is the attention mechanism [3], which enables models to selectively focus on relevant parts of the input, leading to powerful architectures such as Transformers. These models are built from layers combining multi-head self-attention and feed-forward networks, supported by positional encodings to retain word order. Multi-head attention allows the model to capture diverse relationships across the input by processing information in parallel subspaces. With residual connections and layer normalization, Transformers effectively model long-range dependencies and now form the backbone of Large Language Models (LLMs).
@@
As Transformers process increasingly longer sequences, the attention mechanism emerges as a major computational bottleneck [4]. Handling extended contexts in standard attention raises quadratically the number of operations and memory use. To address this, sparse, linear, and low-rank attention [5, 6, 7] have been proposed, which aim to reduce complexity by either simplifying the attention matrix or narrowing focus to the most informative parts of the input. These approaches seek a trade-off between maintaining model accuracy and improving efficiency.
Hardware accelerators for attention focus on managing data locality by keeping queries, keys, and values, in on-chip SRAM [8, 9, 10, 11]. When on-chip storage is insufficient to hold intermediate results across full sequences, data must be written back to external memory, creating performance bottlenecks. These external memory accesses can be reduced by compressing the Transformer’s weights [12], however, to completely overcome this sequence-length dependency, designs such as [13, 14, 15, 16] enhance efficiency by optimizing core attention operations including matrix multiplications and softmax evaluations. Beyond optimizing data movement and computation order, other approaches leverage token similarity [17, 18, 19] to avoid redundant operations, reducing both latency and energy use. Other strategies, like in-memory computation [20], aim to streamline attention processing by performing computation directly within memory arrays.
FlashAttention [21, 22, 23], initially designed for GPUs, has become a highly effective method for speeding up attention computations. It achieves this by combining softmax evaluation with matrix operations in a tiled fashion, enabling efficient parallelism and minimizing memory access overhead. This integrated approach not only shortens execution time but also scales well with longer sequences. As a result, FlashAttention overcomes key performance bottlenecks without sacrificing accuracy of attention outputs.
In this work, we first argue that the simplicity of FlashAttention is not only beneficial for GPUs but also enables the design of efficient parallel hardware accelerators for attention. Second, we move part of FlashAttention’s computations into the logarithmic domain, significantly reducing hardware cost by (a) converting vector-wise multiplications and divisions into additions and subtractions, and (b) avoiding explicit exponential evaluations by leveraging the fused exponential and multiplication operations inherent to the FlashAttention algorithm. For the latter case, a similar approach, but with a narrower scope, has been proposed in [24]. The key advantages of the proposed hardware architecture, named H-FA, are summarized as follows:
@@
  • The reduced dynamic range of logarithmic representation allows the use of fixed-point arithmetic, significantly reducing hardware cost compared to floating-point implementations. Most importantly, the applied logarithmic quantization maintains accuracy for representative LLM applications, as demonstrated by inference on Phi-3.5-mini-instruct [25], Llama-3.2 [26] and Qwen2 [27] models, using various benchmarks for LLM evaluation.
@@ II-A Attention Kernel
In practice, attention scores are scaled by $\sqrt{d}$, where $d$ is the head dimension size, prior to applying softmax normalization [3]. Additionally, some attention scores are masked (i.e., set to a large negative value) to prevent attending to certain positions, such as padding tokens or future tokens in causal attention [3]. To simplify the algorithmic derivation, we omit these features from the following descriptions.
@@
Rather than computing the exponential sum and performing normalization before combining with the value vectors, lazy softmax architectures [13, 17] take a different approach. They simultaneously accumulate both the weighted sum of the value vectors and the sum of the exponentials, deferring the division step until the very end of the process.
@@
Alg. 1 illustrates how attention is computed using the lazy softmax division approach [13, 17]. Initially, it calculates the dot product between the query vector and each key vector. During this step, the maximum score is also identified, which is essential for applying a numerically stable softmax. Next, the output vector $\vec{o}_{i}$ is built incrementally by summing the product of each value vector and its corresponding exponentiated score adjusted by subtracting the maximum score. Simultaneously, the cumulative sum of these exponentials, $\ell_{i}$, is tracked. To obtain the final attention vector for the query $\vec{q}$, the accumulated output is divided by the total sum of exponentials $\ell_{N}$.
@@ II-B FlashAttention
FlashAttention [21], building on the principles of online softmax computation [28], reformulates attention as a streaming process. The enhanced version, FlashAttention-2 [22] that also delays softmax division until the final stage [13, 23], is shown in Alg. 2. Unlike the standard lazy softmax method shown in Alg. 1, FlashAttention-2 performs all intermediate computations including score calculation, exponentiation, and accumulation, within a single inner loop. This avoids the separate step of finding the global maximum score, making it particularly well-suited for handling long input sequences, where efficiency becomes increasingly critical.
@@ III-A Parallel FlashAttention units
FlashAttention-2 avoids the requirement for a full softmax hardware module by decoupling the exponentiation and final division steps, allowing each to be computed independently. Various hardware-friendly techniques have been explored to implement these non-linear operations efficiently. These include piecewise linear approximations applied after range reduction [29], methods based on logarithmic quantization [30], and alternative approximations [31] that express exponentials as powers of two, making them compatible with simple shift-and-add logic that leads to fast and low-cost hardware.
@@ III-B FlashAttention Units operating in parallel to multiple KV sub-blocks
The parallelism available for computing attention can be further exploited for a single query vector by adopting 2D attention techniques [22, 32, 33]. In this approach, the attention output for a given query vector can be computed concurrently across different blocks of the key-value matrices, as long as the intermediate results from each block are appropriately scaled and combined in an online manner.
Specifically, consider the inner loop of Alg. 2, excluding the final normalization step in line 8. Suppose this loop is executed in parallel across two blocks of the $K$ and $V$ matrices, denoted as $(K_{A},V_{A})$ and $(K_{B},V_{B})$. These blocks represent the two halves of the full $K$ and $V$ matrices, respectively. For the same query vector $\vec{q}$, each block independently computes a partial attention output. To reconstruct the final output vector $\vec{o}_{N}$, the partial results $(\vec{o}_{A},\ell_{A})$ and $(\vec{o}_{B},\ell_{B})$, produced by the two halves, must be merged using the following equations, as described in [33, 34]:
@@ IV-A Sum of two products in Logarithm Number System
is computed in the logarithmic number system (LNS) [35]. This formulation will be used in the next section to compute $\ell_{i}$ and $\vec{o}_{i}$ that appear in lines 5 and 6 of Alg. 2.
@@
According to [36], based on the signs of $A$ and $B$, i.e., $s_{A}=s_{xz}$ and $s_{B}=s_{yu}$, we decide how to compute the LNS representation of $c$:
@@ V-A Logarithmic scaled output accumulation
First we remove in (14a) the need to compute logarithm by following Mitchell’s approximation [37]. Since, by definition $0<2^{-|A-B|}<1$, Mitchell has shown in [37] that $\log_{2}(1\pm 2^{-|A-B|})\approx 1\pm 2^{-|A-B|}$. Therefore, $\log_{2}|O_{i}|$ in (14a) can be equivalently written as:
@@ V-A2 Compute $B$
According to (14c) this step involves adding the logarithm of the value vector $V_{i}$ to the quantized $(s_{i}-m_{i})\log_{2}e$. The fixed-point logarithm of the floating-point value vector $V_{i}$ is computed using the method proposed in [38] and further examined in [39].
@@
The approximation $\log_{2}(1+M_{V})\approx M_{V}$ is based on Mitchell’s method [37], which holds since $M$ is a fractional value between 0 and 1.
Exponent $E_{V}$ is an integer and mantissa $M_{V}$ contains only fractional bits. Thus, there is no need to explicitly compute their sum $E_{V}+M_{V}$ that appears in (18). This can be performed implicitly [38] by considering $E_{V}$ and $M_{V}$ as parts of the fixed point number $E_{V}.M_{V}$. The bits dedicated to the integer and fraction part correspond to the original bit width of $E_{V}$ and $M_{V}$. For instance, assuming that $V_{ik}$ follows BFloat16 representation, $E_{V}.M_{V}$ would have 8 integer and 7 fraction bits, respectively. To account for negative logarithm values, using signed (two’s complement) arithmetic, that arise when $V_{ik}<1$, we add one extra bit to the integer part resulting in total of 9 integer bits.
@@
To compute $2^{-f}$ efficiently in hardware, we employ a piecewise-linear (PWL) approximation that exploits the fact that $f$, being a fractional value, is always confined to the range $[0,1)$. This interval is divided into 8 uniform segments using a PWL fitting tool [40], which minimizes the approximation error within each segment. The resulting set of linear coefficients is stored in lookup tables (LUTs), which are indexed using the most significant bits of the fractional input $f$.
@@ VI-A Verification of LLM accuracy
Since part of the computation of FlashAttention is done in the logarithmic domain, it effectively introduces quantization and approximation in multiple arithmetic operations. Our goal is quantify how much these approximations affect the performance of real LLM applications that use FlashAttention kernel. To do so, in the first set of experiments, we emulated the hybrid attention computation kernel ’H-FA’ in Python and integrated it to Microsoft’s Phi-3.5-mini-instruct model. Then, we executed inference tasks on the MMLU benchmark suite [41] using various prompts and utilities provided by the lm-evaluation-harness framework [42]. We simply run inference using the default weights available on the HuggingFace repository [43] as is and without applying any fine-tuning or re-training.
@@
In 18 out the 57 LLM applications the accuracy achieved by both approaches is exactly the same. In some cases, the ’H-FA’ outperforms its full-precision counterpart. This phenomenon is a known artifact of lower-precision models [44]. While it does not indicate that the model is inherently superior, it affirms the fact that the approximations used do not degrade performance. Even in cases that ’H-FA’ has less accuracy the accuracy difference is small and below 5% in the majority of cases. Only in ’global facts’ and ’college chemistry’ accuracy degrades more.
@@
Inference performance was evaluated in a similar way for other contemporary benchmark LLM applications and models. Specifically, we tested Llama3.2-1B and Qwen2-0.5B to analyze how different model sizes perform across four benchmarks: SWAG [45], XCOPA [46], GPQA [47] and GSM8K [48].
Table II presents the accuracy achieved by each LLM on every benchmark. Since each benchmark consists of multiple sub-tasks, as shown in detail for MMLU in Table I, we report the mean accuracy score for each case in Table II. Across all evaluations, the largest observed error is 4%, occurring with the Phi3.5-4B model on GSM8K and SWAG. In all other cases, the error remains below 3%. These results demonstrate that the robustness of our H-FA hybrid computation method is not limited to a specific LLM or benchmark, but also extends to weaker models that are sensitive to numerical inaccuracies, across diverse benchmarks as reported in [49].
@@ VI-C Evaluation of hardware complexity
We implemented both designs in C++ and used Siemens EDA Catapult HLS to synthesize them into Verilog, targeting a 28-nm standard-cell technology. To ensure functional correctness, the C++ implementations were integrated into llama2.c [50], where they produced identical responses to the original implementation across all tested queries. Both accelerators were constrained to operate at a clock frequency of 500 MHz with identical pipelined latency. The total latency varies with the hidden dimension size, requiring 19, 20, and 21 cycles for d=32,64,128d={32,64,128}, respectively. For computing the dot products of a query and a key vector, multi-operand floating-point addition follows the architecture presented in [51].
Physical synthesis of Verilog RTL was carried out using Cadence’s digital implementation tools, and power estimates of the datapaths under comparison were obtained using the Siemens EDA PowerPro analysis and optimization suite. The reported power reflects average consumption measured during inference on various benchmarks using the lm-evaluation-harness framework. The area and power contribution of SRAM memory for implementing KV buffers is also included in the evaluation. For the attention accelerator we assume that it supports a maximum sequence length of $N=1024$ tokens. Thus, each key and value matrix consists of 1024 rows that are distributed to four blocks of 256 rows each. Each row contains $d$ BFloat16 elements (following the size $d$ of the hidden dimension). The area and power of KV SRAM buffers were estimated using Cacti [52] through Accelergy’s hardware components Python library hwcomponents [53] at 2222nm technology node. Acquired results were up-scaled to the 2828nm node by utilizing the DeepScale tool [54].
@@
Platform Process Area Freq. Power Precision Throughput Energy Eff. Area Eff. (nm) (mm2) (MHz) (W) (TOPs or TFLOPs) (TOPs/W) (TOPs/mm2) Keller et al. [9] ASIC 5 0.153 152 - INT4/INT8 3.6/1.8 91.1/39.1 23.53/11.67 MECLA [11] ASIC 28 22.02 1000 2.87 INT8 14 7.08 0.64 FACT [19] ASIC 28 6.03 500 0.337 INT8 1.02 4.39 0.17 Kim et al. [12] ASIC 28 20.25 50 - INT8 3.41 22.9 0.17 Moon et al. [15] ASIC 28 7.29 20 0.002-0.237 AQ 1-8B 0.52 8.94 0.07 Chen et al. [16] ASIC 28 0.636 500 0.108 MXINT4/INT8 0.256 2.37 0.40 COSA plus [14] FPGA 16 - 200 30.3 INT8 1.44 0.05 - TSAcc [18] ASIC 28 8.6 500 3.1 FP32 2.05 0.66 0.24 HFA-1-4 ASIC 28 1.14 500 0.22 Hybrid 0.256(BF16)& 5.41 1.02 (4 KV blocks) BF16&FIX16 0.910(FIX16) HFA-4-4 ASIC 28 3.34 500 0.62 Hybrid 1.64(BF16)& 7.48 1.40 (4 parallel $\vec{q}$, 4 blocks) BF16&FIX16 5.84(FIX16)

### v2

@@ I Introduction
Modern ML and AI systems rely on deep learning models with billions of parameters, reaching human-level performance in tasks like image recognition [vit] and natural language processing [deepseek]. Central to this progress is the attention mechanism [base_attn], which enables models to selectively focus on relevant parts of the input, leading to powerful architectures such as Transformers. These models are built from layers combining multi-head self-attention and feed-forward networks, supported by positional encodings to retain word order. Multi-head attention allows the model to capture diverse relationships across the input by processing information in parallel subspaces. With residual connections and layer normalization, Transformers effectively model long-range dependencies and now form the backbone of Large Language Models (LLMs).
@@
As Transformers process increasingly longer sequences, the attention mechanism emerges as a major computational bottleneck [longformer]. Handling extended contexts in standard attention raises quadratically the number of operations and memory use. To address this, sparse, linear, and low-rank attention [sparse_attn, lin_attn, low_rank_attn] have been proposed, which aim to reduce complexity by either simplifying the attention matrix or narrowing focus to the most informative parts of the input. These approaches seek a trade-off between maintaining model accuracy and improving efficiency.
Hardware accelerators for attention focus on managing data locality by keeping queries, keys, and values, in on-chip SRAM [a3, keller, lu, mecla]. When on-chip storage is insufficient to hold intermediate results across full sequences, data must be written back to external memory, creating performance bottlenecks. These external memory accesses can be reduced by compressing the Transformer’s weights [kim], however, to completely overcome this sequence-length dependency, designs such as [lazy_softmax, cosa, moon, chen] enhance efficiency by optimizing core attention operations including matrix multiplications and softmax evaluations. Beyond optimizing data movement and computation order, other approaches leverage token similarity [elsa, tsacc, fact] to avoid redundant operations, reducing both latency and energy use. Other strategies, like in-memory computation [xformer], aim to streamline attention processing by performing computation directly within memory arrays.
FlashAttention [fa, fa2, nsquared], initially designed for GPUs, has become a highly effective method for speeding up attention computations. It achieves this by combining softmax evaluation with matrix operations in a tiled fashion, enabling efficient parallelism and minimizing memory access overhead. This integrated approach not only shortens execution time but also scales well with longer sequences. As a result, FlashAttention overcomes key performance bottlenecks without sacrificing accuracy of attention outputs.
In this work, we first argue that the simplicity of FlashAttention is not only beneficial for GPUs but also enables the design of efficient parallel hardware accelerators for attention. Second, we move part of FlashAttention’s computations into the logarithmic domain, significantly reducing hardware cost by (a) converting vector-wise multiplications and divisions into additions and subtractions, and (b) avoiding explicit exponential evaluations by leveraging the fused exponential and multiplication operations inherent to the FlashAttention algorithm. For the latter case, a similar approach, but with a narrower scope, has been proposed in [isvlsi-25]. The key advantages of the proposed hardware architecture, named H-FA, are summarized as follows:
@@
  • The reduced dynamic range of logarithmic representation allows the use of fixed-point arithmetic, significantly reducing hardware cost compared to floating-point implementations. Most importantly, the applied logarithmic quantization maintains accuracy for representative LLM applications, as demonstrated by inference on Phi-3.5-mini-instruct [phi], Llama-3.2 [llama3] and Qwen2 [qwen2] models, using various benchmarks for LLM evaluation.
@@ II-A Attention Kernel
In practice, attention scores are scaled by $\sqrt{d}$, where $d$ is the head dimension size, prior to applying softmax normalization [base_attn]. Additionally, some attention scores are masked (i.e., set to a large negative value) to prevent attending to certain positions, such as padding tokens or future tokens in causal attention [base_attn]. To simplify the algorithmic derivation, we omit these features from the following descriptions.
@@
Rather than computing the exponential sum and performing normalization before combining with the value vectors, lazy softmax architectures [lazy_softmax, elsa] take a different approach. They simultaneously accumulate both the weighted sum of the value vectors and the sum of the exponentials, deferring the division step until the very end of the process.
@@
Alg. 1 illustrates how attention is computed using the lazy softmax division approach [lazy_softmax, elsa]. Initially, it calculates the dot product between the query vector and each key vector. During this step, the maximum score is also identified, which is essential for applying a numerically stable softmax. Next, the output vector $\vec{o}_{i}$ is built incrementally by summing the product of each value vector and its corresponding exponentiated score adjusted by subtracting the maximum score. Simultaneously, the cumulative sum of these exponentials, $\ell_{i}$, is tracked. To obtain the final attention vector for the query $\vec{q}$, the accumulated output is divided by the total sum of exponentials $\ell_{N}$.
@@ II-B FlashAttention
FlashAttention [fa], building on the principles of online softmax computation [online-softmax], reformulates attention as a streaming process. The enhanced version, FlashAttention-2 [fa2] that also delays softmax division until the final stage [lazy_softmax, nsquared], is shown in Alg. 2. Unlike the standard lazy softmax method shown in Alg. 1, FlashAttention-2 performs all intermediate computations including score calculation, exponentiation, and accumulation, within a single inner loop. This avoids the separate step of finding the global maximum score, making it particularly well-suited for handling long input sequences, where efficiency becomes increasingly critical.
@@ III-A Parallel FlashAttention units
FlashAttention-2 avoids the requirement for a full softmax hardware module by decoupling the exponentiation and final division steps, allowing each to be computed independently. Various hardware-friendly techniques have been explored to implement these non-linear operations efficiently. These include piecewise linear approximations applied after range reduction [koca], methods based on logarithmic quantization [sole], and alternative approximations [exp2two] that express exponentials as powers of two, making them compatible with simple shift-and-add logic that leads to fast and low-cost hardware.
@@ III-B FlashAttention Units operating in parallel to multiple KV sub-blocks
The parallelism available for computing attention can be further exploited for a single query vector by adopting 2D attention techniques [fa2, block, attention2d]. In this approach, the attention output for a given query vector can be computed concurrently across different blocks of the key-value matrices, as long as the intermediate results from each block are appropriately scaled and combined in an online manner.
Specifically, consider the inner loop of Alg. 2, excluding the final normalization step in line 8. Suppose this loop is executed in parallel across two blocks of the $K$ and $V$ matrices, denoted as $(K_{A},V_{A})$ and $(K_{B},V_{B})$. These blocks represent the two halves of the full $K$ and $V$ matrices, respectively. For the same query vector $\vec{q}$, each block independently computes a partial attention output. To reconstruct the final output vector $\vec{o}_{N}$, the partial results $(\vec{o}_{A},\ell_{A})$ and $(\vec{o}_{B},\ell_{B})$, produced by the two halves, must be merged using the following equations, as described in [attention2d, lean-attn]:
@@ IV-A Sum of two products in Logarithm Number System
is computed in the logarithmic number system (LNS) [log-cnn]. This formulation will be used in the next section to compute $\ell_{i}$ and $\vec{o}_{i}$ that appear in lines 5 and 6 of Alg. 2.
@@
According to [lns-add], based on the signs of $A$ and $B$, i.e., $s_{A}=s_{xz}$ and $s_{B}=s_{yu}$, we decide how to compute the LNS representation of $c$:
@@ V-A Logarithmic scaled output accumulation
First we remove in (14a) the need to compute logarithm by following Mitchell’s approximation [mitchell]. Since, by definition $0<2^{-|A-B|}<1$, Mitchell has shown in [mitchell] that $\log_{2}(1\pm 2^{-|A-B|})\approx 1\pm 2^{-|A-B|}$. Therefore, $\log_{2}|O_{i}|$ in (14a) can be equivalently written as:
@@ V-A2 Compute $B$
According to (14c) this step involves adding the logarithm of the value vector $V_{i}$ to the quantized $(s_{i}-m_{i})\log_{2}e$. The fixed-point logarithm of the floating-point value vector $V_{i}$ is computed using the method proposed in [blinn] and further examined in [sweden].
@@
The approximation $\log_{2}(1+M_{V})\approx M_{V}$ is based on Mitchell’s method [mitchell], which holds since $M$ is a fractional value between 0 and 1.
Exponent $E_{V}$ is an integer and mantissa $M_{V}$ contains only fractional bits. Thus, there is no need to explicitly compute their sum $E_{V}+M_{V}$ that appears in (18). This can be performed implicitly [blinn] by considering $E_{V}$ and $M_{V}$ as parts of the fixed point number $E_{V}.M_{V}$. The bits dedicated to the integer and fraction part correspond to the original bit width of $E_{V}$ and $M_{V}$. For instance, assuming that $V_{ik}$ follows BFloat16 representation, $E_{V}.M_{V}$ would have 8 integer and 7 fraction bits, respectively. To account for negative logarithm values, using signed (two’s complement) arithmetic, that arise when $V_{ik}<1$, we add one extra bit to the integer part resulting in total of 9 integer bits.
@@
To compute $2^{-f}$ efficiently in hardware, we employ a piecewise-linear (PWL) approximation that exploits the fact that $f$, being a fractional value, is always confined to the range $[0,1)$. This interval is divided into 8 uniform segments using a PWL fitting tool [pwlf], which minimizes the approximation error within each segment. The resulting set of linear coefficients is stored in lookup tables (LUTs), which are indexed using the most significant bits of the fractional input $f$.
@@ VI-A Verification of LLM accuracy
Since part of the computation of FlashAttention is done in the logarithmic domain, it effectively introduces quantization and approximation in multiple arithmetic operations. Our goal is quantify how much these approximations affect the performance of real LLM applications that use FlashAttention kernel. To do so, in the first set of experiments, we emulated the hybrid attention computation kernel ’H-FA’ in Python and integrated it to Microsoft’s Phi-3.5-mini-instruct model. Then, we executed inference tasks on the MMLU benchmark suite [mmlu] using various prompts and utilities provided by the lm-evaluation-harness framework [lm-eval]. We simply run inference using the default weights available on the HuggingFace repository [hf] as is and without applying any fine-tuning or re-training.
@@
In 18 out the 57 LLM applications the accuracy achieved by both approaches is exactly the same. In some cases, the ’H-FA’ outperforms its full-precision counterpart. This phenomenon is a known artifact of lower-precision models [i-bert]. While it does not indicate that the model is inherently superior, it affirms the fact that the approximations used do not degrade performance. Even in cases that ’H-FA’ has less accuracy the accuracy difference is small and below 5% in the majority of cases. Only in ’global facts’ and ’college chemistry’ accuracy degrades more.
@@
Inference performance was evaluated in a similar way for other contemporary benchmark LLM applications and models. Specifically, we tested Llama3.2-1B and Qwen2-0.5B to analyze how different model sizes perform across four benchmarks: SWAG [swag], XCOPA [xcopa], GPQA [gpqa] and GSM8K [gsm8k].
Table II presents the accuracy achieved by each LLM on every benchmark. Since each benchmark consists of multiple sub-tasks, as shown in detail for MMLU in Table I, we report the mean accuracy score for each case in Table II. Across all evaluations, the largest observed error is 4%, occurring with the Phi3.5-4B model on GSM8K and SWAG. In all other cases, the error remains below 3%. These results demonstrate that the robustness of our H-FA hybrid computation method is not limited to a specific LLM or benchmark, but also extends to weaker models that are sensitive to numerical inaccuracies, across diverse benchmarks as reported in [weakllms].
@@ VI-C Evaluation of hardware complexity
We implemented both designs in C++ and used Siemens EDA Catapult HLS to synthesize them into Verilog, targeting a 28-nm standard-cell technology. To ensure functional correctness, the C++ implementations were integrated into llama2.c [llama-repo], where they produced identical responses to the original implementation across all tested queries. Both accelerators were constrained to operate at a clock frequency of 500 MHz with identical pipelined latency. The total latency varies with the hidden dimension size, requiring 19, 20, and 21 cycles for d=32,64,128d={32,64,128}, respectively. For computing the dot products of a query and a key vector, multi-operand floating-point addition follows the architecture presented in [ol-align].
Physical synthesis of Verilog RTL was carried out using Cadence’s digital implementation tools, and power estimates of the datapaths under comparison were obtained using the Siemens EDA PowerPro analysis and optimization suite. The reported power reflects average consumption measured during inference on various benchmarks using the lm-evaluation-harness framework. The area and power contribution of SRAM memory for implementing KV buffers is also included in the evaluation. For the attention accelerator we assume that it supports a maximum sequence length of $N=1024$ tokens. Thus, each key and value matrix consists of 1024 rows that are distributed to four blocks of 256 rows each. Each row contains $d$ BFloat16 elements (following the size $d$ of the hidden dimension). The area and power of KV SRAM buffers were estimated using Cacti [cacti] through Accelergy’s hardware components Python library hwcomponents [accelergy] at 2222nm technology node. Acquired results were up-scaled to the 2828nm node by utilizing the DeepScale tool [deepscale].
@@
Platform Process Area Freq. Power Precision Throughput Energy Eff. Area Eff. (nm) (mm2) (MHz) (W) (TOPs or TFLOPs) (TOPs/W) (TOPs/mm2) Keller et al. [keller] ASIC 5 0.153 152 - INT4/INT8 3.6/1.8 91.1/39.1 23.53/11.67 MECLA [mecla] ASIC 28 22.02 1000 2.87 INT8 14 7.08 0.64 FACT [fact] ASIC 28 6.03 500 0.337 INT8 1.02 4.39 0.17 Kim et al. [kim] ASIC 28 20.25 50 - INT8 3.41 22.9 0.17 Moon et al. [moon] ASIC 28 7.29 20 0.002-0.237 AQ 1-8B 0.52 8.94 0.07 Chen et al. [chen] ASIC 28 0.636 500 0.108 MXINT4/INT8 0.256 2.37 0.40 COSA plus [cosa] FPGA 16 - 200 30.3 INT8 1.44 0.05 - TSAcc [tsacc] ASIC 28 8.6 500 3.1 FP32 2.05 0.66 0.24 HFA-1-4 ASIC 28 1.14 500 0.22 Hybrid 0.256(BF16)& 5.41 1.02 (4 KV blocks) BF16&FIX16 0.910(FIX16) HFA-4-4 ASIC 28 3.34 500 0.62 Hybrid 1.64(BF16)& 7.48 1.40 (4 parallel $\vec{q}$, 4 blocks) BF16&FIX16 5.84(FIX16)